OpenStudy (anonymous):

Given f(x) = 4x2 + 3x - 1 and g(x) = 6x + 2, find the function (f + g)(x)

OpenStudy (anonymous):

(f+g)(x)=f(x)+g(x) so just add them and u get, 4x^2+9x+1.

OpenStudy (anonymous):

come to chat brackett

OpenStudy (anonymous):

thanks brackett

OpenStudy (anonymous):

(f+g)(x) = 4xX^2 + 3X - 1 + 6X + 2 = 4X^2 + 9X +1

OpenStudy (anonymous):

thanks acshyamraj look at my last one please!

OpenStudy (anonymous):

YES YOU ARE RIGHT. CONGRATS ! SAMEPHONE18

OpenStudy (anonymous):

Given f(x) = 5x - 6 and g(x) = x2 - 5x + 6, find the function (f + g)(x). Use the ^ key for the exponent. Enter your answer as the example: 4x^3

OpenStudy (anonymous):

x^2

OpenStudy (anonymous):

bracted is right

OpenStudy (anonymous):

brackett I meant, sorry

OpenStudy (anonymous):

:D error is human boss :D